d. 10,000 (For 23-25) You are buying a house for si 50,000 with a 20% d purchase price with a 30 year CPM with bi-weekly payments at 6.125% annual rat own p ayment, the lender will finance the remainder of the e. ( hint: there are 52 weeks in one year) 23. ximately how many payments does it take to reduce the loan balance to $100,0007 a. 268 b. 355 С. 440 d 518 24 What is the total interest payment during the second year of the loan? a. 6,342 b. 6,724 c. 7,217 d. 8,012 25 If at the end of year 5 you made one payment of $20,000 in addition to the regular bi-weekly payment, approximately how many payments will it take thereafter to pay off the loan if you keep making the same payment? a. 362 b. 383 С. 402 d. 438 (For 26-30) A borrower is faced with choosing between two mortgages: Loan B 80,000 25 yr 10.5% Loan amount Term Interest rate Discount points Prepayment penalty 80,000 25 yr 10% 5 pt 0% pt 4%

May I please have help with these three please? also how would I input this in Financial calculator?
Answer #1

Question 25.

Down payment = 0.2*150000 = 30,000

Loan amount = 150,000 -130,000 = 120,000

Number of bi weekly payments = 52/2 = 21 in a year

Number of bi weekly payments in 30 years = 21*30 = 630 payments

Bi weekly payment amount is calculated as =PMT(rate,nper,pv) =PMT(0.06125/21,630,120000) = $416.49

We have the loan amoortization table as shown below:

Bi weekly # Payment Interest Pirncipal Oustanding
0 120000
1 $ 416.49 350 $    66.49 $ 1,19,933.51
2 $ 416.49 349.8061 $    66.68 $ 1,19,866.83
3 $ 416.49 349.6116 $    66.88 $ 1,19,799.95
4 $ 416.49 349.4165 $    67.07 $ 1,19,732.88
5 $ 416.49 349.2209 $    67.27 $ 1,19,665.61
6 $ 416.49 349.0247 $    67.46 $ 1,19,598.15
7 $ 416.49 348.8279 $    67.66 $ 1,19,530.49
8 $ 416.49 348.6306 $    67.86 $ 1,19,462.63
9 $ 416.49 348.4327 $    68.06 $ 1,19,394.57
10 $ 416.49 348.2342 $    68.25 $ 1,19,326.32
11 $ 416.49 348.0351 $    68.45 $ 1,19,257.86
12 $ 416.49 347.8354 $    68.65 $ 1,19,189.21
13 $ 416.49 347.6352 $    68.85 $ 1,19,120.36
14 $ 416.49 347.4344 $    69.05 $ 1,19,051.30
15 $ 416.49 347.233 $    69.26 $ 1,18,982.05
16 $ 416.49 347.031 $    69.46 $ 1,18,912.59
17 $ 416.49 346.8284 $    69.66 $ 1,18,842.93
18 $ 416.49 346.6252 $    69.86 $ 1,18,773.06
19 $ 416.49 346.4214 $    70.07 $ 1,18,703.00
20 $ 416.49 346.2171 $    70.27 $ 1,18,632.72
21 $ 416.49 346.0121 $    70.48 $ 1,18,562.25
22 $ 416.49 345.8066 $    70.68 $ 1,18,491.57
23 $ 416.49 345.6004 $    70.89 $ 1,18,420.68
24 $ 416.49 345.3936 $    71.10 $ 1,18,349.58
25 $ 416.49 345.1863 $    71.30 $ 1,18,278.28
26 $ 416.49 344.9783 $    71.51 $ 1,18,206.77
27 $ 416.49 344.7697 $    71.72 $ 1,18,135.05
28 $ 416.49 344.5606 $    71.93 $ 1,18,063.12
29 $ 416.49 344.3508 $    72.14 $ 1,17,990.98
30 $ 416.49 344.1404 $    72.35 $ 1,17,918.63
31 $ 416.49 343.9294 $    72.56 $ 1,17,846.08
32 $ 416.49 343.7177 $    72.77 $ 1,17,773.30
33 $ 416.49 343.5055 $    72.98 $ 1,17,700.32
34 $ 416.49 343.2926 $    73.20 $ 1,17,627.12
35 $ 416.49 343.0791 $    73.41 $ 1,17,553.71
36 $ 416.49 342.865 $    73.62 $ 1,17,480.09
37 $ 416.49 342.6503 $    73.84 $ 1,17,406.25
38 $ 416.49 342.4349 $    74.05 $ 1,17,332.20
39 $ 416.49 342.2189 $    74.27 $ 1,17,257.93
40 $ 416.49 342.0023 $    74.49 $ 1,17,183.44
41 $ 416.49 341.785 $    74.70 $ 1,17,108.74
42 $ 416.49 341.5672 $    74.92 $ 1,17,033.82
43 $ 416.49 341.3486 $    75.14 $ 1,16,958.68
44 $ 416.49 341.1295 $    75.36 $ 1,16,883.32
45 $ 416.49 340.9097 $    75.58 $ 1,16,807.74
46 $ 416.49 340.6892 $    75.80 $ 1,16,731.94
47 $ 416.49 340.4682 $    76.02 $ 1,16,655.92
48 $ 416.49 340.2464 $    76.24 $ 1,16,579.67
49 $ 416.49 340.0241 $    76.46 $ 1,16,503.21
50 $ 416.49 339.801 $    76.69 $ 1,16,426.52
51 $ 416.49 339.5774 $    76.91 $ 1,16,349.61
52 $ 416.49 339.353 $    77.14 $ 1,16,272.48
53 $ 416.49 339.1281 $    77.36 $ 1,16,195.11
54 $ 416.49 338.9024 $    77.59 $ 1,16,117.53
55 $ 416.49 338.6761 $    77.81 $ 1,16,039.72
56 $ 416.49 338.4492 $    78.04 $ 1,15,961.68
57 $ 416.49 338.2216 $    78.27 $ 1,15,883.41
58 $ 416.49 337.9933 $    78.50 $ 1,15,804.91
59 $ 416.49 337.7643 $    78.72 $ 1,15,726.19
60 $ 416.49 337.5347 $    78.95 $ 1,15,647.23
88 $ 416.49 330.8264 $    85.66 $ 1,13,340.52
89 $ 416.49 330.5765 $    85.91 $ 1,13,254.61
90 $ 416.49 330.3259 $    86.16 $ 1,13,168.45
91 $ 416.49 330.0746 $    86.41 $ 1,13,082.03
92 $ 416.49 329.8226 $    86.67 $ 1,12,995.37
93 $ 416.49 329.5698 $    86.92 $ 1,12,908.45
94 $ 416.49 329.3163 $    87.17 $ 1,12,821.27
95 $ 416.49 329.0621 $    87.43 $ 1,12,733.85
96 $ 416.49 328.8071 $    87.68 $ 1,12,646.17
97 $ 416.49 328.5513 $    87.94 $ 1,12,558.23
98 $ 416.49 328.2948 $    88.19 $ 1,12,470.03
99 $ 416.49 328.0376 $    88.45 $ 1,12,381.58
100 $ 416.49 327.7796 $    88.71 $ 1,12,292.87
101 $ 416.49 327.5209 $    88.97 $ 1,12,203.91
202 $ 416.49 297.0946 $ 119.39 $ 1,01,741.62
203 $ 416.49 296.7464 $ 119.74 $ 1,01,621.87
204 $ 416.49 296.3971 $ 120.09 $ 1,01,501.78
205 $ 416.49 296.0469 $ 120.44 $ 1,01,381.34
206 $ 416.49 295.6956 $ 120.79 $ 1,01,260.55
207 $ 416.49 295.3433 $ 121.15 $ 1,01,139.40
208 $ 416.49 294.9899 $ 121.50 $ 1,01,017.90
209 $ 416.49 294.6355 $ 121.85 $ 1,00,896.05
210 $ 416.49 294.2801 $ 122.21 $ 1,00,773.84
211 $ 416.49 293.9237 $ 122.57 $ 1,00,651.28
212 $ 416.49 293.5662 $ 122.92 $ 1,00,528.35
213 $ 416.49 293.2077 $ 123.28 $ 1,00,405.07
214 $ 416.49 292.8481 $ 123.64 $ 1,00,281.43
215 $ 416.49 292.4875 $ 124.00 $ 1,00,157.43
216 $ 416.49 292.1258 $ 124.36 $ 1,00,033.07
217 $ 416.49 291.7631 $ 124.73 $     99,908.34

So, it takes 216 payments to reduce the loan to 100,000

So nearest answer is 268 weeks (option a)

26. The second year of payments start from payments 22 and last till payment 42. So, the total interest for this from the above table is 7,217

So, answer is $7,217 (option c)

Note, we have answered two MCQ's. Please note that only one MCQ can be answered at a time. Keeping this in mid, we have answered two. So, please post the other separately for experts to answer